# Asymptotic confidence bands for the histogram regression estimator

### v1

@@ 2 Histogram estimator and confidence bands
$\hat{\sigma}^{2}(x)$ $=$ $\frac{\sum_{i=1}^{n}Y_{i}^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}-\hat{m}(x)^{2}$
$=$ $\frac{\sum_{i=1}^{n}(Y_{i}-\hat{m}(x))^{2}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
Then the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}(((\log n)/(n\delta^{p}))^{1/2})$ holds as in Corollary 2.2 under the same assumptions for observations $(X,Y^{2})$, in particular moment conditions and Hölder continuity of $\sigma^{2}(\cdot)$. The histogram estimator $\hat{\sigma}^{2}(\cdot)$ is constant on each hypercube $A_{\delta}(x)$ and thus
@@
is approximated by $\int_{A_{\delta}(x)}\hat{\sigma}^{2}(x)f_{X}(t)\,dt=\hat{\sigma}^{2}(x)p_{\delta}(x)$. Hence, we can estimate $\tau_{\delta}(x)=p_{\delta}(x)^{2}/s_{\delta}(x)$ by
$\hat{\tau}_{\delta}(x)=\frac{\hat{p}_{\delta}(x)}{\hat{\sigma}^{2}(x)}$
@@
The considered classical histogram estimator is a relatively simple estimator, but it has some similarities to random forests, in particular, the piecewise constant approximation. In a current work we apply a similar proof strategy as explained in the next section to derive confidence bands for centered purely random forests. Presumably, this proof strategy can also be used to obtain confidence bands for other nonparametric regression estimators.
@@ 3 Proof strategy
In this section we outline the proof. The estimator error can be decomposed as
$\hat{m}(x)-m(x)=\hat{m}^{(m)}(x)-m(x)+\hat{m}^{(\varepsilon)}(x)$ (11)
with approximation error $\hat{m}^{(m)}(x)-m(x)$ and stochastic error term $\hat{m}^{(\varepsilon)}(x)$, where
$\hat{m}^{(m)}(x)$ $=$ $\sum_{j=1}^{n}m(X_{j})\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\qquad\text{and}$
$\hat{m}^{(\varepsilon)}(x)$ $=$ $\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}.$
First, we discuss a uniform bound on the approximation error. Note that $\hat{m}^{(m)}(x)=\mathbb{E}\left[\hat{m}(x)\mid(X_{j})_{j=1}^{n}\right]$ corresponds to the regression estimate in a setting without observation noise. Therefore, the approximation error is indeed caused by the ability of the piecewise constant estimator to approximate $m$. Moreover, note that the expectation of the approximation error is the bias of the estimator. For any $A\subset\mathbb{R}^{p}$ let us denote its diameter
$\mathfrak{d}(A):=\sup_{x_{1},x_{2}\in A}\|x_{1}-x_{2}\|,$
where $\|\cdot\|$ is the Euclidean norm. Suppose that there exists at least one observation in $A_{\delta}(x)$. Then we apply that $m$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$ with Hölder constant $C_{H}$ to obtain uniformly in $x\in[0,1]^{p}$ that
$|\hat{m}^{(m)}(x)-m(x)|$ $=\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}$
@@
$\leq C_{H}\mathfrak{d}(A_{\delta}(x))^{\alpha}=C_{H}(\sqrt{p}\delta)^{\alpha},$ (12)
where we used $\mathfrak{d}(A_{\delta}(x))=\delta\sqrt{p}$ by the Euclidean geometry of the cells. In particular, the approximation error profits from a smaller diameter or mesh size $\delta$, respectively, of the grid cells.
To handle the stochastic error, we can replace the denominator in $\hat{m}^{(\varepsilon)}(x)$ by its expectation, such that
$\tilde{m}^{(\varepsilon)}(x):=\frac{1}{n}\sum_{j=1}^{n}\varepsilon_{j}p_{\delta}(x)^{-1}\mathbb{I}\{X_{j}\in A_{\delta}(x)\}$ (13)
becomes the asymptotically leading term of the stochastic error $\hat{m}^{(\varepsilon)}(x)$. Handling the remainder term $\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)$ uniformly in $x$ relies on a moment bound for the binomial distribution that can be found in Section 4.3.
The idea for the leading term is that $\tilde{m}^{(\varepsilon)}$ is close to a Gaussian process in $x$ if $n$ is large enough. We write $\tau_{\delta}(x)=p_{\delta}(x)^{2}/s_{\delta}(x)$, see (5) and (10). Note that $p_{\delta}(x)$, $\tau_{\delta}(x)$ and $s_{\delta}(x)$ only depend on the hypercube $A_{\delta}(x)$, so they take only finitely many values for $x\in[0,1]^{p}$. We now calculate the covariance
$\mathrm{Cov}(\tilde{m}^{(\varepsilon)}(x_{1}),\tilde{m}^{(\varepsilon)}(x_{2}))$ $=$ $\frac{1}{n}\mathbb{E}\Big{[}\mathbb{E}\left[\varepsilon^{2}|X\right]p_{\delta}(x_{1})^{-1}p_{\delta}(x_{2})^{-1}\mathbb{I}\{X\in A_{\delta}(x_{1})\}\mathbb{I}\{X\in A_{\delta}(x_{2})\}\Big{]}$
@@
$(n\tau_{\delta}(x))^{1/2}\tilde{m}^{(\varepsilon)}(x)=\mathbb{G}_{n}f_{x,\delta}$ (14)
@@
$\mathcal{F}_{\delta}:=\big{\{}f_{x,\delta}\colon[0,1]^{p}\times\mathbb{R}\to\mathbb{R}\mid f_{x,\delta}(t,\epsilon)=\epsilon\,s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\},x\in[0,1]^{p}\big{\}}.$ (15)
We apply the result by Chernozhukov et al., 2014b formulated in Theorem A.1 in the Appendix A to approximate $\sup_{f\in\mathcal{F}_{\delta}}|\mathbb{G}_{n}f|$ by a sequence of random variables $\mathbf{S}_{\delta}$ with
@@
In particular, this implies $\mathrm{Var}(B_{\delta}(f))=1$ and that the process on the finite function class is in fact a family of independent Gaussian random variables. The application of Theorem A.1 thus leads to the following theorem, which is proved in Section 4.
@@
Let $\log\delta^{-1}=\mathcal{O}(\log n)$. Assume that $\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$ and $\mathbb{E}[|\varepsilon_{1}|^{\nu}]<\infty$ for some $\nu\in[4,\infty)$. Then there exists a sequence of random variables
$\mathbf{S}_{\delta}\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|$
@@
$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right).$
@@
$\mathrm{Var}\left(\tilde{m}^{(\varepsilon)}(x)\right)=\frac{1}{n\tau_{\delta}(x)}=\mathcal{O}\Big{(}\frac{1}{n\delta^{p}}\Big{)},$
@@
i.e. the stochastic error profits from a larger $\delta$, which is caused by larger cells leading to more $\varepsilon_{j}$ being averaged and reducing the variance. Compared to the bound of the approximation error in (12), we recover the typical bias-variance trade-off in nonparametric statistics. The assumptions on $\delta$ in Theorem 2.1 lead to the commonly used undersmoothing. That means $\delta$ has to be chosen such that the stochastic error is slightly larger than the approximation error allowing to build the confidence bands based on the distribution of the stochastic error.
@@ 4 Proofs
Subsequently we gather all the proofs starting with our main result Theorem 2.1. We will write $A\lesssim B$ equivalently for $A=\mathcal{O}(B)$. Moreover, we set $K:=\sup_{x\in[0,1]^{p}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$, and use the notation $\|\cdot\|_{\infty}$ for the supremum norm on $0,1]^{p}$.
@@
We use the decomposition of the estimator and error from (11) and the definition of $\tilde{m}^{(\varepsilon)}$ in (13). For the sequence of random variables $\mathbf{S}_{\delta}$ as defined in Theorem 3.1 we obtain by the reverse triangle inequality that
$\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}$ $\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\mathbf{S}_{\delta}\bigg{|}\leq T_{1}+T_{2}+T_{3},\qquad\text{with}$
$T_{1}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\hat{m}^{(m)}(x)-m(x)\big{|},$ (17)
$T_{2}$ $=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|},$ (18)
$T_{3}$ $=\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}\big{|}\tilde{m}^{(\varepsilon)}(x)-\hat{m}^{(\varepsilon)}(x)|.$ (19)
@@
$\hat{m}^{(m)}(x)-m(x)$ $=\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}$
@@
$\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\Big{|}\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}}\Big{|}\right]$ $\leq C_{H}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathfrak{d}(A_{\delta}(x))^{\alpha}\right]$
$=C_{H}\delta^{\alpha}p^{\alpha/2}.$ (21)
Since the number of cells is equal to $\delta^{-p}$, we obtain
$\mathbb{E}$ $\left[\sup_{x\in[0,1]^{p}}|m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1)|\right]$
$\leq\|m\|_{\infty}\mathbb{E}\left[\sup_{x\in[0,1]^{p}}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\right]$
$\leq\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}\mathbb{E}\big{[}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\delta}(x)\}\big{]}$
$=\|m\|_{\infty}\delta^{-p}\sup_{x\in[0,1]^{p}}(1-p_{\delta}(x))^{n}\leq\|m\|_{\infty}\delta^{-p}(1-c_{X}\delta^{p})^{n}.$
To handle (17), note that $n^{1/2}\sup_{x}|\tau_{\delta}(x)|^{1/2}\leq(C_{X}/c_{\sigma^{2}})^{1/2}n^{1/2}\delta^{p/2}$ by (8). Together with (20) and (21) we obain for all $\kappa>0$
$\mathbb{P}\left(T_{1}\geq\frac{\kappa}{\log n}\right)$ $=\mathbb{P}\left(\|m-\hat{m}^{(m)}\|_{\infty}\geq\kappa\frac{(c_{\sigma^{2}}/C_{X})^{1/2}}{n^{1/2}\delta^{p/2}\log n}\right)$
$\lesssim\mathbb{E}\big{[}\|m-\hat{m}^{(m)}\|_{\infty}\big{]}\kappa^{-1}(n\delta^{p})^{1/2}\log n$
$\lesssim(\delta^{2\alpha}n\delta^{p})^{1/2}\log n+\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n.$
The first term is $o(1)$ with (3) and for the second term we have
$\delta^{-p}(1-c_{X}\delta^{p})^{n}(n\delta^{p})^{1/2}\log n\lesssim\exp(-c_{X}n\delta^{p})\delta^{-p}(n\delta^{p})^{1/2}\log n\to 0$
because $\delta\leq n^{-1/(2\alpha+p)}$ by (3) and $\delta^{p}n\geq n^{2\nu}$ by (4). Therefore, $T_{1}=o_{\mathbb{P}}((\log n)^{-1})$.

### v2

@@ 2 Histogram estimator and confidence bands
$\widehat{\sigma}_{\Delta}^{2}(x)$ $=$ $\frac{\sum_{i=1}^{n}Y_{i}^{2}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}-\widehat{m}(x)^{2}$
$=$ $\frac{\sum_{i=1}^{n}(Y_{i}-\widehat{m}(x))^{2}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}.$
Then the uniform rate of convergence $\mathcal{O}_{\mathbb{P}}(\sqrt{(\log n)\Delta/n})$ holds as in Corollary 2.5 under the same assumptions for observations $(X,Y^{2})$, in particular under moment conditions and Hölder continuity of $\sigma^{2}(\cdot)$. The histogram estimator $\widehat{\sigma}_{\Delta}^{2}(\cdot)$ is constant on each subset $A_{\Delta}(x)$ and thus
@@
is approximated by $\int_{A_{\Delta}(x)}\widehat{\sigma}_{\Delta}^{2}(x)f_{X}(t)\,dt=\widehat{\sigma}_{\Delta}^{2}(x)p_{\Delta}(x)$. Hence, we can estimate $\tau_{\Delta}(x)=p_{\Delta}(x)^{2}/s_{\Delta}(x)$ by
$\widehat{\tau}_{\Delta}(x)=\frac{\widehat{p}_{\Delta}(x)}{\widehat{\sigma}_{\Delta}^{2}(x)}$
@@ 3 Proof strategy
In this section we outline the proof. The estimation error can be decomposed as
$\widehat{m}(x)-m(x)=\widehat{m}^{(m)}(x)-m(x)+\widehat{m}^{(\varepsilon)}(x)$ (15)
with approximation error $\widehat{m}^{(m)}(x)-m(x)$ and stochastic error $\widehat{m}^{(\varepsilon)}(x)$, where
$\widehat{m}^{(m)}(x)$ $=$ $\sum_{j=1}^{n}m(X_{j})\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}\qquad\text{and}$
$\widehat{m}^{(\varepsilon)}(x)$ $=$ $\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}.$
First, we discuss a uniform bound on the approximation error. Note that $\widehat{m}^{(m)}(x)=\mathbb{E}\left[\widehat{m}(x)\mid(X_{j})_{j=1}^{n}\right]$ corresponds to the regression estimate in a setting without observation noise. Therefore, the approximation error is indeed caused by the ability of the piecewise constant estimator to approximate $m$. Moreover, note that the expectation of the approximation error is the bias of the estimator. Suppose that there exists at least one observation in $A_{\Delta}(x)$. Then we use that $m$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$ with Hölder constant $C_{H}$ to obtain uniformly in $x\in\mathcal{X}$ that
$|\widehat{m}^{(m)}(x)-m(x)|$ $=\Big|\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}\Big|$
@@
$\leq C_{H}\operatorname{diam}(A_{\Delta}(x))^{\alpha}=C_{H}C_{p}^{\alpha}\Delta^{-\alpha/p},$ (16)
where we used $\operatorname{diam}(A_{\Delta}(x))\leq C_{p}\Delta^{-1/p}$ by Assumption 2.1. In particular, the approximation error profits from a finer partition.
To handle the stochastic error, we can replace the denominator in $\widehat{m}^{(\varepsilon)}(x)$ by its expectation, such that
$\widetilde{m}^{(\varepsilon)}(x)=\frac{1}{n}\sum_{j=1}^{n}\varepsilon_{j}\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{p_{\Delta}(x)}$ (17)
becomes the asymptotically leading term of the stochastic error $\widehat{m}^{(\varepsilon)}(x)$. Handling the remainder term $\widehat{m}^{(\varepsilon)}(x)-\widetilde{m}^{(\varepsilon)}(x)$ uniformly in $x$ relies on a moment bound for the binomial distribution that can be found in Section 6.3.
The idea for the leading term is that $\widetilde{m}^{(\varepsilon)}$ is close to a Gaussian process in $x$ if $n$ is large enough. We write $\tau_{\Delta}(x)=p_{\Delta}(x)^{2}/s_{\Delta}(x)$, see (8) and (14). Note that $p_{\Delta}(x)$, $\tau_{\Delta}(x)$ and $s_{\Delta}(x)$ only depend on $x$ via the cell $A_{\Delta}(x)$, so they take only finitely many values for $x\in\mathcal{X}$. We now calculate the covariance
$\mathrm{Cov}(\widetilde{m}^{(\varepsilon)}(x_{1}),\widetilde{m}^{(\varepsilon)}(x_{2}))$ $=\frac{1}{n}\mathbb{E}\Big[\mathbb{E}\left[\varepsilon^{2}|X\right]\frac{\mathbb{I}\{X\in A_{\Delta}(x_{1})\}\mathbb{I}\{X\in A_{\Delta}(x_{2})\}}{p_{\Delta}(x_{1})p_{\Delta}(x_{2})}\Big]$
@@
$(n\tau_{\Delta}(x))^{1/2}\widetilde{m}^{(\varepsilon)}(x)=\mathbb{G}_{n}f_{x,\Delta}$ (18)
@@
$\mathcal{F}_{\Delta}:=\big\{f_{x,\Delta}\colon\mathcal{X}\times\mathbb{R}\to\mathbb{R}\mid f_{x,\Delta}(t,\epsilon)=\epsilon\,s_{\Delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\Delta}(x)\},x\in\mathcal{X}\big\}.$ (19)
We apply the result by Chernozhukov et al., 2014b formulated in Theorem B.1 in the Appendix B to approximate $\sup_{f\in\mathcal{F}_{\Delta}}|\mathbb{G}_{n}f|$ by a sequence of random variables $\mathbf{S}_{\Delta}$ with
@@
In particular, this implies $\mathrm{Var}(B_{\Delta}(f_{x,\Delta}))=1$ and that the process on the finite function class is in fact a family of independent Gaussian random variables. The application of Theorem B.1 thus leads to the following theorem, which is proved in Section 6.
@@
Grant Assumption 2.1 and let $\log\Delta=\mathcal{O}(\log n)$. Assume that $\mathbb{E}[|\varepsilon_{1}|^{\nu}]<\infty$ for some $\nu\in[4,\infty)$ and $\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$. Then there exists a sequence of random variables
$\mathbf{S}_{\Delta}\overset{d}{=}\max_{j=1,\ldots,\Delta}|Z_{j}|$
@@
$\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}\Delta^{1/2}}{n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}\Delta^{1/4}}{n^{1/4}}+\frac{(\log n)\Delta^{1/6}}{n^{1/6}}\right).$
@@
$\mathrm{Var}\left(\widetilde{m}^{(\varepsilon)}(x)\right)=\frac{1}{n\tau_{\Delta}(x)}=\mathcal{O}\Big(\frac{\Delta}{n}\Big),$
@@
i.e. the stochastic error profits from a partition into fewer cells as larger cells lead to more $\varepsilon_{j}$ being averaged. Compared to the bound of the approximation error in (16), we recover the typical bias-variance trade-off in nonparametric statistics. The assumptions on $\Delta$ in Theorem 2.2 lead to the commonly used undersmoothing. That means $\Delta$ has to be chosen such that the stochastic error is slightly larger than the approximation error, allowing the confidence bands to be built based on the distribution of the stochastic error.
4 Numerical illustration
To illustrate our theoretical performance in a simulation example, we consider the following model: Let $p=2$, $X\sim U([0,1]^{2})$ and the regression function by given by
$m(x)=3\sqrt{\big|0.36-\|x\|^{2}\big|}-(x_{1}+x_{2}),\qquad x=(x_{1},x_{2})\in[0,1]^{2}.$
Note that $m$ is Hölder regular of order $\alpha\leq 1/2$ and that $m$ is close the a radial function. We choose homogeneous Gaussian noise with known noise level $\sigma>0$ and fix the sample size $n=800$. We use histogram estimators based on two different partitions. First, we use a histogram regression estimator $\widehat{m}_{\text{square}}$ based on a partition into $\Delta=h^{-2}$ squares of edge length $h=1/6$. Second, we consider an estimator $\widehat{m}_{\text{circ}}$ which relies on a partition based on circular segments. More precisely, we decompose the unit square into 15 circles and split the middle ones into two to four segments. Both partitions consist of $\Delta=36$ cells. For comparison we also calculate a Nadaraya-Watson estimator $\widehat{m}_{\text{nw}}$ based on a Gaussian kernel. The bandwidth is chosen either as $h/\sqrt{12}$, such that the standard deviation of the kernel corresponds to that of a uniform distribution on a $h\times h$ square, or via the automatic bandwidth choice implemented in the R package np. The true regression function, both histogram estimators as well as the Nadaraya-Watson estimator are illustrated in Figure 2.
Figure 2: Regression function (top left) together with realizations of the Nadaraya-Watson estimator with automatic bandwidth choice (top right) and the two histogram estimators (bottom row) with two different partitions based on $n=800$ observations with noise level $\sigma=1$.
$\sigma$ 1 2 3 4 5
$R(\widehat{m}_{\text{circ}})$ 0.9010 1.4501 2.0383 2.6376 3.2394
$R(\widehat{m}_{\text{square}})$ 1.1895 1.5800 2.0947 2.6338 3.1805
$R(\widehat{m}_{\text{nw}})$ (fixed $h$) 1.0076 1.8232 2.7005 3.5900 4.4815
$R(\widehat{m}_{\text{nw}})$ (adaptive $h$) 0.9446 1.2055 1.4417 1.6066 1.7530
CP for $\widehat{m}_{\text{circ}}$ $c_{\Delta}(0.95)$ 0.99 1.00 1.00 1.00 1.00
$\widetilde{c}_{\Delta}(0.95)$ 1.00 1.00 1.00 1.00 1.00
CP for $\widehat{m}_{\text{square}}$ $c_{\Delta}(0.95)$ 0.98 0.99 0.99 1.00 1.00
$\widetilde{c}_{\Delta}(0.95)$ 0.98 1.00 1.00 1.00 1.00
UCP for $\widehat{m}_{\text{circ}}$ $c_{\Delta}(0.95)$ 0.05 0.61 0.78 0.85 0.88
$\widetilde{c}_{\Delta}(0.95)$ 0.15 0.71 0.85 0.93 0.95
UCP for $\widehat{m}_{\text{square}}$ $c_{\Delta}(0.95)$ 0.00 0.15 0.49 0.58 0.61
$\widetilde{c}_{\Delta}(0.95)$ 0.00 0.26 0.54 0.63 0.70
Table 1: Mean uniform estimation errors R(m^)R($\widehat${m}) for histogram estimators based on circular and quadratic partitions and mean uniform estimation error of a Nadaraya-Watson estimator; average pointwise coverage probabilities (CP) and uniform coverage probabilities (UCP) of the confidence bands with nominal level $0.95$ based on both histogram estimators with exact and asymptotic critical value, all for different noise levels $\sigma$.
We compare the performance of the three estimators in terms of the mean uniform estimation error $R(\widehat{m})=\mathbb{E}[\sup_{x\in[0,1]^{2}}|\widehat{m}(x)-m(x)|]$ based on a grid on $0,1]^{2}$ with mesh size 0.005 and a Monte Carlo approximation of the expectation with 100 Monte Carlo iterations. Moreover, we calculate the uniform empirical coverage probabilities of the confidence bands with nominal confidence level $0.95$ based on the two histogram estimators and using the critical values $c_{\Delta}$ and $\widetilde{c}_{\Delta}$ from (7) and (12), respectively. As an additional feedback on the coverage properties of the confidence bands, we also report the average pointwise coverage probabilities over the spatial grid. Since the radius of the confidence bands based on $\widetilde{c}_{\Delta}$ are by the factor $\frac{\widetilde{c}_{16}(0.95)}{c_{16}(0.95)}\approx 1.0453$ larger than those based on $c_{\Delta}$, the coverage probabilities of the latter are throughout smaller. To see the dependence of the signal-to-noise ratio we vary the noise level $\sigma\in\{1,2,\dots,5\}$. All results are summarized in Table 1. We see that for small noise levels the performance of the histogram estimators is comparable to that of the Nadaraya-Watson estimator, while the coverage of the confidence bands is poor. By increasing the noise level, we enter an under-smoothing regime such that the estimation error becomes dominated by the stochastic error and the coverage considerably improves. Naturally, the estimation error grows with $\sigma$. The increasing deviation to the (adaptive) Nadaraya-Watson estimator emphasizes that the partitions are coarse in these cases.
The partition of $\widehat{m}_{\text{circ}}$ is better adapted to the geometry of the regression function. As a consequence, the coverage probabilities for $\widehat{m}_{\text{circ}}$ are considerably better than for $\widehat{m}_{\text{square}}$ in all cases. Moreover, for noise levels $\sigma=1,2,3$ the estimation error of $\widehat{m}_{\text{circ}}$ is smaller, while for $\sigma=4,5$ the estimation error of $\widehat{m}_{\text{square}}$ is only slightly better.
5 Conclusion
Histogram regression estimators are used a lot in applications and serve as a building block for more advanced statistical methods such as random forests. Aiming for uniform statistical inference, we construct histogram based confidence bands which allow for a general choice of partitions. This flexibility can improve the finite sample behaviour of the histogram estimator depending on the data structure.
Our confidence band is applicable for possibly irregular nonparametric regression functions with Hölder-regularity only of order $\alpha\in(0,1]$. The authors are not aware of any uniform confidence band for non-differentiable regression functions in the literature. These mild model assumptions are possible due to a proof structure, which deviates from more classical proof methods and which could also be transferred to other methods and models. For instance, the current preprint Neumeyer et al., (2025) demonstrates that a similar proof strategy can be used to derive confidence bands for centered purely random forests. Some ideas to obtain confidence bands for other nonparametric regression estimators are discussed in Appendix A for different choices of cells including the case of a Nadaraya-Watson estimator.
@@ 6 Proofs
Subsequently we gather all the proofs starting with our main result Theorem 2.2. We will write $A\lesssim B$ equivalently for $A=\mathcal{O}(B)$. Moreover, we set $K=\sup_{x\in\mathcal{X}}\mathbb{E}[|\varepsilon|^{3}|X=x]<\infty$, and use the notation $\|\cdot\|_{\infty}$ for the supremum norm on $\mathcal{X}$.
@@
We use the decomposition of the estimator and error from (15) and the definition of $\widetilde{m}^{(\varepsilon)}$ in (17). For the sequence of random variables $\mathbf{S}_{\Delta}$ as defined in Theorem 3.1 we obtain by the reverse triangle inequality that
$\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}$ $\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|-\mathbf{S}_{\Delta}\bigg|\leq T_{1}+T_{2}+T_{3},\qquad\text{with}$
$T_{1}$ $=\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}\big|\widehat{m}^{(m)}(x)-m(x)\big|,$ (21)
$T_{2}$ $=\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|,$ (22)
$T_{3}$ $=\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}\big|\widetilde{m}^{(\varepsilon)}(x)-\widehat{m}^{(\varepsilon)}(x)|.$ (23)
@@
$\widehat{m}^{(m)}(x)-m(x)$ $=\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}$
@@
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}\Big|\sum_{j=1}^{n}(m(X_{j})-m(x))\frac{\mathbb{I}\{X_{j}\in A_{\Delta}(x)\}}{\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}}\Big|\right]$ $\leq C_{H}\mathbb{E}\left[\sup_{x\in\mathcal{X}}\operatorname{diam}(A_{\Delta}(x))^{\alpha}\right]$
$=C_{H}C_{p}^{\alpha}\Delta^{-\alpha/p}.$ (25)
Moreover, we obtain by (10)
$\mathbb{E}$ $\left[\sup_{x\in\mathcal{X}}|m(x)(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\Delta}(x)\}-1)|\right]$
$\leq\|m\|_{\infty}\mathbb{E}\left[\sup_{x\in\mathcal{X}}\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\Delta}(x)\}\right]$
$\leq\|m\|_{\infty}\Delta\sup_{x\in\mathcal{X}}\mathbb{E}\big[\mathbb{I}\{\nexists j\in[n]:X_{j}\in A_{\Delta}(x)\}\big]$
$=\|m\|_{\infty}\Delta\sup_{x\in\mathcal{X}}(1-p_{\Delta}(x))^{n}\leq\|m\|_{\infty}\Delta(1-c_{X}{c_{p}\Delta^{-1}})^{n}.$
To handle (21), note that $n^{1/2}\sup_{x}|\tau_{\Delta}(x)|^{1/2}\leq(C_{X}C_{p}/c_{\sigma^{2}})^{1/2}(n/\Delta)^{1/2}$ by (11). Together with (24) and (25) we obtain for all $\kappa>0$
$\mathbb{P}\left(T_{1}\geq\frac{\kappa}{\log n}\right)$ $=\mathbb{P}\left(\|m-\widehat{m}^{(m)}\|_{\infty}\geq\kappa\frac{(c_{\sigma^{2}}/(C_{X})C_{p})^{1/2}\Delta^{1/2}}{n^{1/2}\log n}\right)$
$\lesssim\mathbb{E}\big[\|m-\widehat{m}^{(m)}\|_{\infty}\big]\kappa^{-1}(n/\Delta)^{1/2}\log n$
$\lesssim{(\Delta^{-2\alpha/p}n/\Delta)^{1/2}\log n+\Delta(1-c_{X}c_{p}\Delta^{-1})^{n}(n/\Delta)^{1/2}\log n}.$
The first term is $o(1)$ due to (5) and for the second term we have
$\Delta(1-c_{X}c_{p}\Delta^{-1})^{n}(n/\Delta)^{1/2}\log n\lesssim\exp(-c_{X}c_{p}n/\Delta)\Delta(n/\Delta)^{1/2}\log n\to 0$
because $\Delta\geq n^{1/(2\alpha/p+1)}$ by (5) and $n/\Delta\geq n^{2\nu}$ by (6). Therefore, $T_{1}=o_{\mathbb{P}}((\log n)^{-1})$.